# ARTEMIS: Autoregressive End-to-End Trajectory Planning with Mixture of Experts for Autonomous Driving

### v1

@@ Abstract
This paper presents ARTEMIS, an end-to-end autonomous driving framework that combines autoregressive trajectory planning with Mixture-of-Experts (MoE). Traditional modular methods suffer from error propagation, while existing end-to-end models typically employ static one-shot inference paradigms that inadequately capture the dynamic changes of the environment. ARTEMIS takes a different method by generating trajectory waypoints sequentially, preserves critical temporal dependencies while dynamically routing scene-specific queries to specialized expert networks. It effectively relieves trajectory quality degradation issues encountered when guidance information is ambiguous, and overcomes the inherent representational limitations of singular network architectures when processing diverse driving scenarios. Additionally, we use a lightweight batch reallocation strategy that significantly improves the training speed of the Mixture-of-Experts model. Through experiments on the NAVSIM dataset, ARTEMIS exhibits superior competitive performance, achieving 87.0 PDMS and 83.1 EPDMS with ResNet-34 backbone, demonstrates state-of-the-art performance on multiple metrics.
@@ I Introduction
Autonomous driving has experienced rapid development over the past few decades. Traditional modular methods compartmentalize autonomous driving tasks into discrete modules such as perception, prediction, and planning[1, 2, 3]. However, cumulative errors and complex interdependencies between these modules can become constrained at predefined interfaces. End-to-end models overcome these issues by mapping raw sensor data directly to planned trajectories or control signals[4, 5, 6, 7]. However, their static, one-shot inference paradigm often fails to capture the dynamic evolution of the environment (Fig.1 (a)). In contrast, autoregressive methods generate trajectories sequentially, preserving temporal coherence and allowing for adaptive decision-making based on previously planned segments. Autoregressive models have been widely applied in the field of trajectory prediction[8, 9, 10]. Recent research has begun to explore unified frameworks that simultaneously accomplish world model construction and complete or partial trajectory planning tasks through autoregressive modeling methods[11, 12]. However, the current problem is that single-network end-to-end models still struggle to adequately capture and adapt to diverse driving scenarios.
@@
To tackle the inherent complexity in trajectory planning, researchers are increasingly adopting sophisticated architectures like the Mixture of Experts (MoE) framework[19]. MoE employs multiple specialized expert networks along with intelligent routing mechanisms to dynamically distribute and process inputs—a strategy that has yielded significant success in large-scale language models[20, 21, 22]. In autonomous driving, the planning trajectories generated by end-to-end models inherently encompass multiple potential behavioral modalities, reflecting the fundamental uncertainty in driving behavior. Drivers may select from several reasonable future actions under identical environmental conditions (Fig.1 (b))[15, 16, 17], and traditional single-network architectures struggle to accurately characterize this intrinsic diversity of behavioral patterns. In contrast, MoE enables expert modules to focus on specific driving scenarios or behavioral patterns, thereby learning characteristic distributions of driving behavior without reliance on predefined guidance signals. This endogenous multimodal modeling method effectively circumvents potential trajectory quality degradation issues that may arise when guidance information deviates from actual conditions. Recently, diffusion-based models (Fig.1 (c))[18], [23] have introduced novel generative modeling paradigms to autonomous driving, demonstrating enhanced trajectory diversity. Although these methodologies have established state-of-the-art performance in end-to-end autonomous driving, they typically employ static paradigms that generate all trajectory points simultaneously (or through multiple denoising iterations), limiting their capacity to accurately capture the dynamic evolutionary characteristics of trajectory development. In comparison, autoregressive methods utilizing MoE demonstrate superior temporal sequence capturing capabilities, environmental adaptability, and practical utility, as they can operate without requiring strong prior constraints.
@@ II-A1 End-to-End Autonomous Driving
End-to-end autonomous driving systems aim to directly map raw sensor data to vehicle trajectories. Formally, we define end-to-end autonomous driving as a conditional sequence generation problem where the input is a sequence of historical sensor observations $S=(s_{1},s_{2},...,s_{T})$, where $T$ denotes the history horizon. The model is required to generate a trajectory sequence for future time instants $Y=(y_{1},y_{2},...,y_{H})$, where $H$ denotes the prediction horizon. Each point $y_{h}\in\mathcal{R}^{d}$, with $d$ representing the dimensionality of each waypoint, which contains information such as position and orientation. The objective of an end-to-end model is to learn the conditional probability distribution $p(Y|S)$.
@@
where $y_{<t}$ denotes all trajectory points generated prior to time $t$. This factorization allows the model to consider the previously generated trajectory when producing each new waypoint, thereby preserving the temporal coherence of the trajectory.
@@ II-A2 Mixture of Experts
MoE is designed for the dynamic handling of complex input data by leveraging multiple expert networks and an intelligent routing mechanism that enables the dynamic allocation of computational resources and network models. In this work, we adopt a design similar to DeepseekMoE [20], where some experts are designated as shared experts to capture generalizable knowledge and reduce the redundancy of routing experts. Given an input $x\in\mathbb{R}^{n\times d_{model}}$, where $n$ denotes the sequence length and $d_{\text{model}}$ represents the feature dimensionality, the output of our MoE can be formalized as:
@@ II-B4 Training Loss
Consistent with some end-to-end autonomous driving methods, we adopt a staged training method to mitigate training instability[4, 5, 15, 27]. Specifically, the perception network and its auxiliary tasks, including semantic mapping and object detection, are first trained. Subsequently, the entire network is trained in an end-to-end method. This training strategy significantly enhances both the stability and the overall performance of the model. Furthermore, we opted not to employ the expert balance loss commonly utilized in MoE architecture, as its application in datasets with imbalanced feature distributions could potentially impede the acquisition of specialized strategic knowledge by individual experts.
@@
End-to-End Training Loss. In this stage, the entire network is trained in an end-to-end method. The overall loss is defined as:
@@
In the “Input” column, “C” denotes the use of camera images as sensor input, and “C&L” represents the use of both camera images and lidar as sensor inputs. All models require ego-vehicle state as input, while “S” indicates that the model only needs ego-vehicle state as input. * For fair comparison, we used the official scores of versions with the same backbone network and hidden layer dimensions. The best and second-best scores are highlighted in bold and underlined, respectively.
@@ III-C Implementation Details
We employ the Transfuser [14] as our perception network, with ResNet34 as the backbone for feature extraction. The perception inputs consist of concatenated images from the front-left, front, and front-right cameras, as well as point cloud data covering a $64\,\text{m}\times 64\,\text{m}$ area. In the autoregressive planning module that incorporates MOE, we configure $E_{private}=5$ domain-specific experts along with $E_{shared}=1$ shared expert. During forward propagation, the top $K=2$ experts with the highest score are selected for activation. The model is trained on the navtrain split using two A100 GPUs with a batch size of 128. The initial learning rate is set to $2\times 10^{-4}$ and the weight decay is set to $1\times 10^{-4}$. The model performs 8 autoregressive steps, with each step outputting one trajectory point comprising the $x$, $y$ and heading. After all autoregressive steps, the generated trajectory is refined to form a 4-second planning trajectory at 2 Hz. During the training stage of the perception module, the loss weight coefficients $\lambda_{\text{sem}}$, $\lambda_{\text{class}}$, and $\lambda_{\text{box}}$ are set to 10, 10, and 5, respectively. In the end-to-end training stage, the weight coefficients $\lambda_{\text{sem}}$, $\lambda_{\text{class}}$, $\lambda_{\text{box}}$, $\lambda_{\text{traj}}$, $\lambda_{\text{var}}$, and $\lambda_{\text{NLL}}$ are set to 2, 2, 1, 15, 0.01, and 0.5, respectively.
@@ III-E Ablation Study
The Routing Network. To verify the effectiveness of the intrinsic routing mechanism, we compare it with an explicit driving‑command guidance scheme. The experimental results in Tab.IV show that the intrinsic router effectively prevents inappropriate assignment of domain‑specific experts when explicit guidance deviates from the actual scene.
@@
Cascaded Refinement Layers. Tab.V reports how varying the number of refinement layers influences model performance. Increasing the cascade depth improves performance up to two layers, beyond this point, the gain stabilises and the marginal benefit diminishes.
Table V: Impact of the Number of Refinement Layers.
@@
Domain‑Specific Experts. Tab.VI investigates the impact of varying the number of experts. Enlarging the expert pool from three to five gradually increases the model’s performance, indicating enhanced capability for handling complex scenarios. However, expanding to ten experts leads to a performance drop of 1.5 points, suggesting that, given limited training data, an excessive number of experts disperses resources and causes functional overlap.
Table VI: Effect of the Number of Domain‑Specific Experts.
@@
Batch Reallocation on Training Speed. To assess the contribution of batch reallocation, we compared training speed with and without this strategy across different batchsize. As presented in Tab.VII, under identical hardware conditions, batch reallocation significantly accelerates training, the training samples per second from 19.2 to 43.5 as the batch size grows from 64 to 256. Although reallocation introduces extra overhead, it is negligible relative to expert network computation and is offset by the gain in parallel efficiency.
Table VII: Effect of Batch Reallocation on Training Samples per Second under Different Batchsizes (Using a Single NVIDIA A100 GPU).

### v2

@@ Abstract
This paper presents ARTEMIS, an end-to-end autonomous driving framework that combines autoregressive trajectory planning with Mixture-of-Experts (MoE). Traditional modular methods suffer from error propagation, while existing end-to-end models typically employ static one-shot inference paradigms that inadequately capture the dynamic changes of the environment. ARTEMIS takes a different method by generating trajectory waypoints sequentially, preserves critical temporal dependencies while dynamically routing scene-specific queries to specialized expert networks. It effectively relieves trajectory quality degradation issues encountered when guidance information is ambiguous, and overcomes the inherent representational limitations of singular network architectures when processing diverse driving scenarios. Additionally, we use a lightweight batch reallocation strategy that significantly improves the training speed of the Mixture-of-Experts model. Through experiments on the NAVSIM dataset, ARTEMIS exhibits superior competitive performance, achieving 87.0 PDMS and 83.1 EPDMS with ResNet-34 backbone, demonstrates state-of-the-art performance on multiple metrics. Code will be available under https://github.com/Lg0914/ARTEMIS.
@@ I Introduction
Autonomous driving has experienced rapid development over the past few decades. Traditional modular methods compartmentalize autonomous driving tasks into discrete modules such as perception, prediction, and planning[1, 2, 3]. However, cumulative errors and complex interdependencies between these modules can become constrained at predefined interfaces. End-to-end models overcome these issues by mapping raw sensor data directly to planned trajectories or control signals[4, 5, 6]. However, their static, one-shot inference paradigm often fails to capture the dynamic evolution of the environment (Fig.1 (a)). In contrast, autoregressive methods generate trajectories sequentially, preserving temporal coherence and allowing for adaptive decision-making based on previously planned segments. Autoregressive models have been widely applied in the field of trajectory prediction[7, 8, 9]. Recent research has begun to explore unified frameworks that simultaneously accomplish world model construction and complete or partial trajectory planning tasks through autoregressive modeling methods[10, 11]. However, the current problem is that single-network end-to-end models still struggle to adequately capture and adapt to diverse driving scenarios.
@@
To tackle the inherent complexity in trajectory planning, researchers are increasingly adopting sophisticated architectures like the Mixture of Experts (MoE) framework[18]. MoE employs multiple specialized expert networks along with intelligent routing mechanisms to dynamically distribute and process inputs—a strategy that has yielded significant success in large-scale language models[19, 20, 21]. In autonomous driving, the planning trajectories generated by end-to-end models inherently encompass multiple potential behavioral modalities, reflecting the fundamental uncertainty in driving behavior. Drivers may select from several reasonable future actions under identical environmental conditions (Fig.1 (b))[14, 15, 16], and traditional single-network architectures struggle to accurately characterize this intrinsic diversity of behavioral patterns. In contrast, MoE enables expert modules to focus on specific driving scenarios or behavioral patterns, thereby learning characteristic distributions of driving behavior without reliance on predefined guidance signals. This endogenous multimodal modeling method effectively circumvents potential trajectory quality degradation issues that may arise when guidance information deviates from actual conditions. Previous work, which was based on structured data representations as input[22], incorporated Mixture-of-Experts (MoE) into planning tasks and demonstrated strong performance on the NuPlan dataset[23]. Recently, diffusion-based models (Fig.1 (c))[17], [24] have introduced novel generative modeling paradigms to autonomous driving, demonstrating trajectory diversity. Although these methodologies have established state-of-the-art performance in end-to-end autonomous driving, they typically employ static paradigms that generate all trajectory points simultaneously (or through multiple denoising iterations), limiting their capacity to accurately capture the dynamic evolutionary characteristics of trajectory development. In comparison, autoregressive methods utilizing MoE demonstrate superior temporal sequence capturing capabilities, environmental adaptability, and practical utility, as they can operate without requiring strong prior constraints.
@@ II-A1 End-to-End Autonomous Driving
Formally, we define end-to-end autonomous driving as a conditional sequence generation problem where the input is a sequence of historical sensor observations $S=(s_{1},s_{2},...,s_{T})$, where $T$ denotes the history horizon. The model is required to generate a trajectory sequence for future time instants $Y=(y_{1},y_{2},...,y_{H})$, where $H$ denotes the prediction horizon. Each point $y_{h}\in\mathcal{R}^{d}$, with $d$ representing the dimensionality of each waypoint. The objective of an end-to-end model is to learn the conditional probability distribution $p(Y|S)$.
@@
where $y_{<t}$ denotes trajectory points generated prior to time $t$.
@@ II-A2 Mixture of Experts
In this work, we adopt a design similar to DeepseekMoE [19], where some experts are designated as shared experts to capture generalizable knowledge and reduce the redundancy of routing experts. Given an input $x$, the output of our MoE can be formalized as:
@@ II-B4 Training Loss
Consistent with some end-to-end methods, we adopt a staged training method to mitigate training instability[4, 5, 14]. Specifically, the perception network and its auxiliary tasks, including semantic mapping and object detection, are first trained. Subsequently, the entire network is trained in an end-to-end method. This training strategy significantly enhances both the stability and the overall performance of the model. Furthermore, we opted not to employ the expert balance loss commonly utilized in MoE architecture, as its application in datasets with imbalanced feature distributions could potentially impede the acquisition of specialized strategic knowledge by individual experts.
@@
End-to-End Training Loss. In this stage, the entire network is trained in end-to-end method. The overall loss is defined as:
@@
In the “Input” column, “C” denotes the use of camera images as sensor input, and “C&L” represents the use of both camera images and lidar as sensor inputs. All models require ego-vehicle state as input. * For fair comparison, we used the official scores of versions with the same backbone network and hidden layer dimensions. The best and second-best scores are highlighted in bold and underlined, respectively.
@@ III-C Implementation Details
We employ the Transfuser [13] as our perception network, with ResNet34 as the backbone for feature extraction. The perception inputs consist of concatenated images from the front-left, front, and front-right cameras, as well as point cloud data covering a $64\,\text{m}\times 64\,\text{m}$ area. In the autoregressive planning module that incorporates MOE, we configure $E_{private}=5$ domain-specific experts along with $E_{shared}=1$ shared expert. During forward propagation, the top $K=2$ experts with the highest score are selected for activation. The model is trained on the navtrain split using two A100 GPUs with a batch size of 128. The initial learning rate is set to $2\times 10^{-4}$ and the weight decay is set to $1\times 10^{-4}$. The model performs 8 autoregressive steps, with each step outputting one trajectory point comprising the $x$, $y$ and heading. After all autoregressive steps, the generated trajectory is refined to form a 4-second planning trajectory at 2 Hz. During the training stage of the perception module, the loss weight coefficients $\lambda_{\text{sem}}$, $\lambda_{\text{class}}$, and $\lambda_{\text{box}}$ are set to 10, 10, and 5, respectively. In the end-to-end training stage, the weight coefficients $\lambda_{\text{sem}}$, $\lambda_{\text{class}}$, $\lambda_{\text{box}}$, $\lambda_{\text{traj}}$, and $\lambda_{\text{NLL}}$ are set to 1, 1, 0.5, 15, and 0.2, respectively.
@@ III-E Ablation Study
The Routing Network. To verify the effectiveness of the intrinsic routing mechanism, we compare it with an explicit driving‑command guidance scheme. The experimental results in Tab.IV show that the intrinsic routing mechanism effectively prevents inappropriate assignment of domain‑specific experts when explicit guidance deviates from the actual scene.
We also conducted experiments by fixing the activation to single experts within the model. The results presented in Table.V demonstrate that activating any individual expert in isolation fails to achieve optimal performance, underscoring the necessity of the routing network for facilitating effective expert allocation. Moreover, the results confirms that the model training process did not exhibit disproportionate preference toward utilizing any particular expert, indicating balanced expert utilization throughout the learning process.
@@
Table V: Performance of Single Expert Activation Strategies
Individually Activated Expert NC DAC EP TTC C PDMS
Expert 1 97.5 92.5 79.0 92.2 99.9 83.4
Expert 2 98.0 94.7 81.0 93.7 100 86.2
Expert 3 97.4 91.0 77.4 91.9 99.8 81.9
Expert 4 97.8 93.8 80.0 93.3 99.9 85.1
Expert 5 98.1 94.1 80.2 93.2 100 85.4
Intrinsic Routing 98.3 95.1 81.4 94.3 100 87.0
Cascaded Refinement Layers. Tab.VI reports how varying the number of refinement layers influences model performance. Increasing the cascade depth improves performance up to two layers, beyond this point, the gain stabilises and the marginal benefit diminishes.
Table VI: Impact of the Number of Refinement Layers.
@@
Domain‑Specific Experts. Tab.VII investigates the impact of varying the number of experts. Enlarging the expert pool from three to five gradually increases the model’s performance, indicating enhanced capability for handling complex scenarios. However, expanding to ten experts leads to a performance drop of 1.5 points, suggesting that, given limited training data, an excessive number of experts disperses resources and causes functional overlap.
Table VII: Effect of the Number of Domain‑Specific Experts.
@@
Batch Reallocation on Training Speed. To assess the contribution of batch reallocation, we compared training speed with and without this strategy across different batchsize. As presented in Tab.VIII, under identical hardware conditions, batch reallocation significantly accelerates training, the training samples per second from 19.2 to 43.5 as the batch size grows from 64 to 256. Although reallocation introduces extra overhead, it is negligible relative to expert network computation and is offset by the gain in parallel efficiency.
Table VIII: Effect of Batch Reallocation on Training Samples per Second under Different Batchsizes (Using a Single NVIDIA A100 GPU).